Daily Summary – May 20th

Crew 181 Daily Summary Report

May 20th, 2017


MDRS Daily Summary Report for SOL 7


Summary Title: Third Quarter Effect in Full Effect

Mission Status: Now that all crew members are accustomed to one another, we are feeling the third-quarter effect and getting antsy for upcoming days.

Sol Activity Summary: After a late night at the observatory, only one crew member woke up in time for the scheduled morning EVA. Therefore, the EVA start time was pushed by one hour. Regardless, it was successful for everyone a part of it. Cassandra made breakfast (French toast) for everyone. Janet and Charlie continued to troubleshoot the observatory (with much success).

With the loss of honorary crew member Olly and the stressful events of this upcoming week, we seem to know each other well enough now to realize when idiosyncrasies are driving us crazy. Just a few days ago, the second floor of the Hab seemed like the place to hang out. Now people have settled in different areas – the first floor, the science lab – to work on their individual projects. We will try to work together to gain back some of our crew collaboration time. I hope this isn’t the first signs of the third quarter effect. We are only to the half way mark!

Afternoon was filled with writing, photo editing, and Settlers of Catan. Evening EVA was great in the sunset light, making the landscape a greater saturation of reds than we thought possible. We are now all settling in, the smell of Janet and Charlie’s dinner permeating the second floor.

Look Ahead Plan: More observatory work is in the talks. We are submitting an EVA request for tomorrow. We are gearing up (slowly) for our Martian shipment of Ozzy and Jack Osbourne on Monday (we may be out of sim for the day and reports may be coming in late.) We are also disheartened by Avishek having to leave the mission early to return back to PhD work on Earth (also on Monday.)

Anomalies In Work: None at the current time.

Weather: Bright, sunny, lower than average temperature.

Crew Physical Status: Everyone is okay.

EVA: One EVA took place around the Hab, the other took place at White Moon.

Reports to be Filed: journalist report, engineering report, EVA report, and EVA request.

Support Requested: None at the current time.


Prepared by Cassandra



Cassandra Klos

Commander, Crew 181